8. The magnitude of the gravitational field on the surface of a particular planet is 2g. The planets mass is half the mass of Earth. What is the planets radius in terms of the radius R of Earth? (B) RE/2 (C) Rel (D) 2RE
Answer #1

g = GM/Re^2

where M = mass of earth , Re = radius of earth

2g = G*(M/2) / (R)^2

4g = GM/R^2

dividind both equations we have

1/4 = (R/Re)^2

R = Re / 2

so option B) Re / 2 is the correct answer